Calculate the volume of 15.98 g of SO2 at 25 degrees C and 1 atm of pressure.
Answer #1

Answer

By using ideal gas equation

PV= nRT

P= pressure = 1 atm   

V= volume = ?

n= number of moles of SO2 = mass/molar mass = 15.98/64.066= 0.249 moles

R= gas constnat = 0.0821 LatmmolK

T = temperature = 25+273.15= 298.15 K

V= nRT/P = (0.249*0.0821*298.15)/1 = 6.1 L